OpenStudy (anonymous):

The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle. A. 18° B. 36° C. 54° D. 58°

jimthompson5910 (jim_thompson5910):

Let x and y be the two complementary angles since they are complementary, we know they add to 90 degrees, so x+y = 90

jimthompson5910 (jim_thompson5910):

we are told that "The sum of the measures of two complementary angles exceeds the difference of their measures by 72", so x+y = x-y + 72 y = -y + 72 y + y = 72 2y = 72 y = ???

OpenStudy (mathstudent55):

Complementary angles have measures that add up to 90. Let one angle have measure x. The other angle has measure 90 - x Sum of measures: x + (90 - x) Difference of their measures: x - (90 - x) Sum of measures exceeds diff of measures by 72: x + (90 + x) = x - (90 - x) + 72 Solve for x

OpenStudy (anonymous):

so whats the answer

OpenStudy (mathstudent55):

x + (90 - x) = x - (90 - x) + 72 x - x + 90 = x - 90 + x + 72 90 = 2x - 18 108 = 2x 54 = x 90 - 54 = 36 The smaller angle measure 36 degrees.